OpenStudy (anonymous):

Find a particular solution to x'=x+2y+2t, y'=3x+2y-4. a)using integrating factor method, b) using eigenvector decomposition, c)using undetermined coeffiecients

OpenStudy (anonymous):

I don't know about the first two, so I'll just stick with the last method. \[\left(\begin{matrix}x\\y\end{matrix}\right)'=\left(\begin{matrix}1&2\\3&2\end{matrix}\right)\left(\begin{matrix}x\\y\end{matrix}\right)+t\left(\begin{matrix}2\\0\end{matrix}\right)+\left(\begin{matrix}0\\-4\end{matrix}\right)\] Find the corresponding characteristic solution using the usual eigenvalue method: \[\left(\begin{matrix}x\\y\end{matrix}\right)'=\left(\begin{matrix}1&2\\3&2\end{matrix}\right)\left(\begin{matrix}x\\y\end{matrix}\right)_c\] As a guess for the particular solution, try \[\left(\begin{matrix}x\\y\end{matrix}\right)_p=t\vec{a}+\vec{b}\] This gives you \[\left(\begin{matrix}x\\y\end{matrix}\right)_p'=\color{red}{\vec{a}= \left( \begin{matrix} 1&2\\ 3&2 \end{matrix} \right) \left(t\vec{a}+\vec{b}\right) +t \left( \begin{matrix} 2\\ 0 \end{matrix} \right) + \left( \begin{matrix} 0\\ -4 \end{matrix} \right)}\] Let \(A=\left(\begin{matrix}1&2\\3&2\end{matrix}\right)\). This gives you \[\vec{a}=At\vec{a}+A\vec{b}+t\left(\begin{matrix}2\\0\end{matrix}\right)+\left(\begin{matrix}0\\-4\end{matrix}\right)\] Matching up "coefficients," you get \[\begin{cases} A\vec{a}+\left(\begin{matrix}2\\0\end{matrix}\right)=\vec{0}\\ A\vec{b}+\left(\begin{matrix}0\\-4\end{matrix}\right)=\vec{a} \end{cases}\] Once you figure out \(\vec{a}\) and \(\vec{b}\), you should get the right solution. I'd check with WolframAlpha just to be sure.